OpenStudy (anonymous):

Complete the identity. sin4x - cos4x = ? A. 1 - 2cos2 x B. 1 - 2sin2 x C. 1 + 2sin2 x D. 1 + 2cos2 x

OpenStudy (anonymous):

@UnkleRhaukus @jim_thompson5910

terenzreignz (terenzreignz):

What do you mean by sin4x... do you mean \[\sin^{4} x\] ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

sin^4x - cos^4x = ? A. 1 - 2cos^2 x B. 1 - 2sin^2 x C. 1 + 2sin^2 x D. 1 + 2cos^2 x

terenzreignz (terenzreignz):

Well, it's A, you want me to elaborate? :)

OpenStudy (anonymous):

yes plz

OpenStudy (anonymous):

\[(\sin ^{2}x-\cos ^{2}x)(\sin ^{2}x+\cos ^{2}x)\] \[(\sin ^{2}x-\cos ^{2}x)(1) \implies (1-\cos ^{2}x-\cos ^{2}x)\implies (1-2\cos ^{2}x)\]

terenzreignz (terenzreignz):

That ^

OpenStudy (anonymous):

the identity used is \[\sin ^{2}x+\cos ^{2}x=1\]

jimthompson5910 (jim_thompson5910):

sin^4x - cos^4x (sin^2x)^2 - (cos^2x)^2 (sin^2x-cos^2x)(sin^2x+cos^2x) (sin^2x-cos^2x)(1) sin^2x-cos^2x (1-cos^2x)-cos^2x 1-cos^2x-cos^2x 1-2cos^2x So sin^4x - cos^4x = 1-2cos^2x for all values of x

OpenStudy (anonymous):

@nitz i dont get this part